The speed of a river current is 4 mph. If a boat travels 20 miles downstream in the same time that it takes to travel 10 miles​ upstream, find the speed of the boat in still water.

Answer by Boreal(15235)

current is 4 mph
speed of boat =x
downstream, speed is 4+x and time is 20/x+4, because time is distance over speed.
upstream, speed is 4-x and time is 10/x-4
The two times are equal
so 20/x+4=10/x-4
20x-80=10x+40
10x=120
x=12 mph
16 mph downstream and takes 1h15m to go 20 miles (1.25 hr)
8 mph upstream and takes 1h15m to go 10 miles.